OpenStudy (mathmath333):

Probabilities that Rajesh passes in Physics, Math and Chemistry are p, m and c respectively. Of these subjects, Rajesh has 75% chance of passing in at least one, 50% chance of passing in at least two and 40% chance of passing in exactly two. Find which of the following is true.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{Probabilities that Rajesh passes in Physics, Math and Chemistry }\hspace{.33em}\\~\\ & \normalsize \text{are p, m and c respectively. Of these subjects, Rajesh has }\hspace{.33em}\\~\\ & \normalsize \text{75% chance of passing in at least one, 50% chance of passing}\hspace{.33em}\\~\\ & \normalsize \text{in at least two and 40% chance of passing in exactly two. Find}\hspace{.33em}\\~\\ & \normalsize \text{which of the following is true.}\hspace{.33em}\\~\\ & a.)\ p+m+c=\dfrac{19}{20} \hspace{.33em}\\~\\ & b.)\ p+m+c=\dfrac{27}{20} \hspace{.33em}\\~\\ & c.)\ pmc=\dfrac{1}{20} \hspace{.33em}\\~\\ & d.)\ pmc=\dfrac18 \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

i found this but cant understand the answer https://math.stackexchange.com/questions/1422953/probability-problem-possibly-based-on-principle-of-inclusion-exclusion

OpenStudy (mathmate):

Yes, principle of inclusion and exclusion is your answer. Let's start with a venn diagram so that we're on the same wavelength. 

OpenStudy (mathmate):